# On admissibility in post-hoc hypothesis testing

### v2

@@ Appendix C On the relationship to Bayesian decision theory
A common response to the drawbacks of classical Neyman-Pearson theory that we’ve highlighted is to pivot to Bayesian decision theory (Bernardo et al., 1994). While the Bayesian often/typically omits the notion of type-I error control and focuses solely on loss minimization, we show below that he gains the ability to handle post-hoc loss functions for free. We also discuss what a Bayesian notion of type-I risk control might look like, and how to achieve it.
@@
A fundamental result is that the Bayes optimal decision rule (Lehmann and Casella, 1998, Theorem 4.1.1), henceforth ‘Bayes decision’ for brevity, is always given by $\delta_{\pi}$, meaning that it satisfies $B_{\pi}(\delta_{\pi})=\inf_{\delta}B_{\pi}(\delta)$ where
@@
We emphasize that $\phi_{\pi}$ is well-defined since, as mentioned earlier, $\phi$ is allowed to see $B(X)$. We also note that this result is not new. For example, Grünwald (2023) notes that that Bayesian decision can accommodate post-hoc loss functions, but we hope the formal treatment above clarifies in what sense this is true, and allows for easier comparison to our frequentist results.
@@ Appendix D On risk control vs error probabilities
The first argument is historical. As we’ve discussed, Wald (1939) noted that standard Neyman-Pearson hypothesis testing is a part of classical statistical decision theory once it is reformulated in terms of loss functions. And decision theory has always been concerned with expected loss as the error metric, due to compelling arguments such as representation theorems (Savage, 1954), coherence and Dutch books arguments (De Finetti, 1937), complete class theorems (Wald, 1950; Blackwell and Girshick, 1954), amenability to central limit theorems and laws of large numbers, and so on. Thus, our emphasis on risk control can be seen as a return to the roots of hypothesis testing as decision theory, where expected risk is the norm.
The second is a pragmatic argument. As we mentioned in the introduction, there are alternatives to Neyman-Pearson hypothesis testing, the most successful of which is the Bayesian agenda. But if one is fond of the tools of the Neyman-Pearson paradigm—such as significance levels, power, and accept/reject procedures—and seeks to generalize those tools to handle post-hoc analysis, then the formulation of post-hoc hypothesis testing studied here (and in Grünwald, 2024) appears to be the only option currently on offer. (See Appendix C for a longer discussion on Bayesian decision theory and how it relates to our framework.)

### v3

@@ Appendix C On the relationship to Bayesian decision theory
A common response to the drawbacks of classical Neyman-Pearson theory that we’ve highlighted is to pivot to Bayesian decision theory (bernardo1994bayesian). While the Bayesian often/typically omits the notion of type-I error control and focuses solely on loss minimization, we show below that he gains the ability to handle post-hoc loss functions for free. We also discuss what a Bayesian notion of type-I risk control might look like, and how to achieve it.
@@
A fundamental result is that the Bayes optimal decision rule (lehmann1998theory, Theorem 4.1.1), henceforth ‘Bayes decision’ for brevity, is always given by $\delta_{\pi}$, meaning that it satisfies $B_{\pi}(\delta_{\pi})=\inf_{\delta}B_{\pi}(\delta)$ where
@@
We emphasize that $\phi_{\pi}$ is well-defined since, as mentioned earlier, $\phi$ is allowed to see $B(X)$. We also note that this result is not new. For example, grunwald2023posterior notes that that Bayesian decision can accommodate post-hoc loss functions, but we hope the formal treatment above clarifies in what sense this is true, and allows for easier comparison to our frequentist results.
@@ Appendix D On risk control vs error probabilities
The first argument is historical. As we’ve discussed, wald1939contributions noted that standard Neyman-Pearson hypothesis testing is a part of classical statistical decision theory once it is reformulated in terms of loss functions. And decision theory has always been concerned with expected loss as the error metric, due to compelling arguments such as representation theorems (savage1954foundations), coherence and Dutch books arguments (de1937prevision), complete class theorems (wald1950statistical; blackwell1954theory), amenability to central limit theorems and laws of large numbers, and so on. Thus, our emphasis on risk control can be seen as a return to the roots of hypothesis testing as decision theory, where expected risk is the norm.
The second is a pragmatic argument. As we mentioned in the introduction, there are alternatives to Neyman-Pearson hypothesis testing, the most successful of which is the Bayesian agenda. But if one is fond of the tools of the Neyman-Pearson paradigm—such as significance levels, power, and accept/reject procedures—and seeks to generalize those tools to handle post-hoc analysis, then the formulation of post-hoc hypothesis testing studied here (and in grunwald2024beyond) appears to be the only option currently on offer. (See Appendix C for a longer discussion on Bayesian decision theory and how it relates to our framework.)